OpenStudy (anonymous):

A certain chemical reaction experiences a 7.50 × 10-7 kilogram loss in mass. If this mass were to be totally converted to energy, what would be the magnitude of that energy? (1 J = 1 kg times m squared over s squared ) 1.20 × 1023 J 2.5 × 109 J 4.00 × 101 J 6.75 × 1010 J @abmon98

OpenStudy (anonymous):

@Abmon98 do you get it I was thinking it was B but I may be wrong ?

OpenStudy (anonymous):

are u there @Abmon98 please I need you help

OpenStudy (abmon98):

@veronica.montanez i am trying but i dont quite get the m^2/s^2 part

OpenStudy (anonymous):

isn't it like e=mc^2 or no

OpenStudy (abmon98):

(m/s)^2=m^2/s^2 oh yes you are right

OpenStudy (anonymous):

so how would I do that

OpenStudy (abmon98):

E=( 7.50 × 10-7)(3.00*10^8)^2

OpenStudy (abmon98):

10^-7**

OpenStudy (anonymous):

is that for the 3.00

OpenStudy (abmon98):

no for the 7.50

OpenStudy (anonymous):

ok its D and I need you help on three more is that ok :) ?

OpenStudy (abmon98):

i got Energy as 67500000000 so its a D

OpenStudy (abmon98):

okay sure but i have to do something now i be right back in 15 minutes .

OpenStudy (anonymous):

ok @Abmon98